solve by substitution 5x-3y= -43
-8x+y=66 what is the solution set I came up with ( -8,3) am I right

Answer by algebrapro18(249)

No that solution isn't right. First off you want to solve the second equation for y.
-8x+y=66
y = 66+8x
Now we substitute that value into the first equation and solve for x.
5x-3y=-43
5x-3(66+8x) = -43
5x-198-24x=-43
-19x-198 = -43
-19x=155
x = -155/19
Now we substitute that into the equation we found for y.
y = 66+8x = 66+8(-155/19) = 14/19.
So the solution to the system is (-155/19, 14/19)